OpenStudy (anonymous):

Solve the following system of equations. x + 3y - z = 2 x - 2y + 3z = 7 x + 2y - 5z = -21 Can anyone explain how to do this to me, please?

OpenStudy (anonymous):

You can do this with a graphing calculator ( cheat) or you can do it on paper.

OpenStudy (anonymous):

How do you do it on paper, like, work it out?

OpenStudy (anonymous):

x-2y+3z=7 (second problem) x+2y-5z=-21(third problem) Combine x-2z=-14

OpenStudy (anonymous):

x+3y-z=2 x-2y+3z=7 you want one of the variables to cancel each other out. 2x+6y-2z=4 (multiply by 2) 3x-6y+9z=21 (multiply by 3) combine 5x+7z=25